OpenStudy (anonymous):

F (x)=3-2x+x^2 find f(-3)

OpenStudy (anonymous):

is f(x)= F(x)? or we have to use integrals?

OpenStudy (whpalmer4):

I'm assuming we are simply supposed to evaluate \(f(x) = 3-2x+x^2\) at \(x = -3\): \[f(-3) = 3 - 2(-3) + (-3)^2 = \]

OpenStudy (anonymous):

0 is the answer

OpenStudy (anonymous):

F(-3)=3+6+9=18

OpenStudy (whpalmer4):

@faisalalif1999 how on earth did you come up with 0 from the expression \[f(-3) = 3 -2(-3) + (-3)^2\]\[(-2)*(-3) = 6\]\[(-3)^2 = (-3)*(-3) = 9\]\[f(-3) = 3-2(-3)+(-3)^2 = 3+6+9 = 18\]

OpenStudy (anonymous):

Reason.. about the question if x = -3.. F(-3) = 3 - 2(-3) + (-3)^2 F(-3) = 3 - (-6) + 9 >> (-3)^2 = 9 // 3 x 3 = 9 // 2(-3) = 6 F(-3) = 3 + 6 + 9 = 18 F(-3) = 18... This would be the answer of F(-3)...

OpenStudy (anonymous):

f = -3

OpenStudy (anonymous):

sorry, x = 18 lol